AEON: The OSM tried to repeat the Rapid Response targets, even after they had been observed sucessfully. I sent an email to Hannah with more details, so that she can check what happened.
AEON: We should add a check for the filter set in NEWFIRM, after the "acquisition done" button is pressed. We had an observation today where we had to check the telescope focus before we started a sequence in HX filter. This means that we switched to JX for improving the focus. When done, we hit the "Acquisition done" button and the NEWFIRM fields were re-populated correctly; except for switching the filter back to HX. Adding this to the system would be good.
2nd Half: At ~5:30 local time, an image had been taken. The Exposure Status section of the GUI had the "busy" LED illuminated after readout and the controls were disabled in grey. We shutdown the NEWFIR GUI and restated. Upon the restart we lost control of the filter wheel and spent some tie (~15-20 min) trying to get the filter wheel to initialize. Upon regaining control of the filter wheel we were able to resume science observations.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2026-01-30 | 8.0 | 0.0 | 0.5 | 0.0 | 0.0 | 8.5 |
AEON observations on the first half of the night:
- We checked the fix by Hannah for the re-population of fields in NEWFIRM after the "acquisition done" button is pressed. The behaviour was as expected.
- Tomás entered a couple of "Rapid Response" targets to LCO as a test, in preparation for the Rubin alerts test. The requests arrived fine and data was taken.